If Einstein E is total energy absorbed by 1 mole of a substance and that each molecule absorbs 1 quantum of energy, then calculate the value of E in terms of λ in cm.

Grade:12

1 Answers

askiitian.expert- chandra sekhar
14 years ago

Hi Ankur,

E1 = hc/λ ------ energy absorbed by one molecule

E = Nhc/λ -------energy absorbed by one mole

N =6.023* 10^23

h= 6.625*10^-34 J-sec

c = 3 * 10^10 cm/sec

λ in cm

E = 3.99/(λ in cm)   J